### 28 Nov 2023 v1

# GeoScaler: Geometry and Rendering-Aware Downsampling of 3D Mesh Textures

Abstract
High-resolution texture maps are necessary for representing real-world objects accurately with 3D meshes. The large sizes of textures can bottleneck the real-time rendering of high-quality virtual 3D scenes on devices having low computational budgets and limited memory. Downsampling the texture maps directly addresses the issue, albeit at the cost of visual fidelity. Traditionally, downsampling of texture maps is performed using methods like bicubic interpolation and the Lanczos algorithm. These methods ignore the geometric layout of the mesh and its UV parametrization and also do not account for the rendering process used to obtain the final visualization that the users will experience. Towards filling these gaps, we introduce GeoScaler, which is a method of downsampling texture maps of 3D meshes while incorporating geometric cues, and by maximizing the visual fidelity of the rendered views of the textured meshes. We show that the textures generated by GeoScaler deliver significantly better quality rendered images compared to those generated by traditional downsampling methods.
1 Introduction
Textured meshes have proven to provide efficient representations of 3D objects in terms of representation density, quantization, and ease of rendering. For example, point clouds are often too sparse, voxel grids suffer from memory inefficiencies, and volumetric representations are difficult to render. Textured meshes are less affected by these limitations, and because of their versatility, they are ubiquitous in real-time computer graphics applications like video games, and most graphical hardware products are designed and manufactured to accelerate their rendering. However, real-time, high-quality graphics applications are still difficult to realize on low-compute devices such as virtual reality headsets and smartphones. This has forced developers and graphic artists to either build meshes containing smaller texture maps or downsample their existing high-resolution texture maps, leading to a compromise in visual quality.
Contemporaneous methods such as bilinear, bicubic, and Lanczos [23] resampling are very commonly used for downsampling textures. These methods do not adapt to the data geometry, as they were designed for planar two-dimensional images. Planar images differ from texture maps in two key ways, making standard resamplers inefficient for processing 3D mesh textures.
The two ways that texture maps differ are that they contain discontinuities and warping artifacts that arise during UV mapping, which is the process of projecting the 3D surface of a mesh onto a 2D plane. As it is often impossible to flatten a 3D surface onto a two-dimensional plane without causing major distortions, it is required to slice the 3D surfaces into a number of patches, then flatten each patch separately. Fig. 1 illustrates the trade-off between these two types of artifacts using different projections of the spherical surface of the Earth onto a 2D map. The Mercator projection (Fig. 1a) of the world map significantly distorts areas near the two poles by expanding their surface areas more than on regions closer to the equator. Conversely, the Goode homolosine projection [12] (Fig. 1b) reduces the distortions significantly but introduces significant discontinuities.
The aforementioned resampling methods operate by applying fixed-size symmetric kernels on images, to estimate the resampled values. Applying simple interpolation kernels at the discontinuities on texture maps leads to blur and/or boundary displacements. Due to the localized nature of the applied kernels, true neighbors of a pixel sitting on a discontinuity in a texture map may not be involved while filtering that pixel. Another problem arises from the symmetric nature of standard interpolation kernels, whereby pixels in all directions are weighted similarly. However, the assumption of symmetric distances is not generally true for texture maps. Textures become warped when the 3D surfaces they are painted on are flattened onto 2D planes, as the UV parametrization is always non-isometric for non-planar surfaces: areas of the triangles on the mesh in 3D are non-uniformly modified when mapped to their 2D representation on the UV plane. Ideally, the downsampling algorithm should account for warping by flexibly weighing the pixels during the interpolation
Figure 1: The trade-off between warping and discontinuities when creating UV parametrization. (a) Regions like Greenland and Antarctica which lie near the poles consume disproportionately large areas on Mercator projection. (b) Goode homolosine projection reduces warping artifacts but introduces many discontinuities.
Furthermore, texture maps are viewed by users in the form of rendered 2D views. Thus, it is crucial to ensure that the downsampling process is designed to optimize the perceived visual quality of the rendered content, instead of prioritizing data fidelity of the texture on the UV plane. Finally, synthesis of the UV parametrization might cause inefficient usage of space on the texture map due to the presence of a large number of discontinuities. This can lead to wasted memory consumption during rendering and may detrimentally impact the visual quality of images rendered using downsampled texture maps.
These observations demonstrate the interdependence between the geometry of a mesh, its UV parametrization, and its texture map. In our view, this motivates the need for the development of downsampling techniques for texture maps that account for these characteristics of texture maps, as well as the rendering process used to generate the final view of a mesh.
Towards this end, we have created GeoScaler, which is a texture map downsampling method that addresses the drawbacks of existing methods. An overview of our method is shown in Fig. 2. GeoScaler is a deep neural network (DNN) based downsampling model that is optimized to maximize the perceptual quality of images rendered using the downsampled texture map of the mesh.
The design of GeoScaler is based on three key concepts:
  1. 1.
    Rendering loss is calculated between views rendered from the mesh with its original texture map, and the mesh with a downsampled texture map generated by the DNN. This loss is used to optimize the network.
  2. 2.
    The GeoCoding module exploits the geometric layout and the UV parametrization of the mesh to counteract the effects of warping distortions and discontinuities in the texture map
  3. 3.
    The UVWarper module computes a warping function that adjusts the UV parameters of all the vertices of the mesh to obtain a more efficient layout of the texture.
In the experiments, we demonstrate that GeoScaler significantly outperforms existing downsampling methods on a wide variety of meshes by providing up to 3dB and 2dB improvement for 8x and 4x downsampling of textures respectively, against rendered views generated from the meshes with the original textures.
Figure 2: Overview of the GeoScaler architecture. For a given input textured 3D mesh, an optimal texture downsampling function is learned which incorporates information of the geometry of the mesh and its UV parametrization. The model parameters are iteratively updated via differentiable rendering by computing the loss between views rendered from the original texture of the mesh versus the downsampled texture.
2 Related Works
2.1 Generation of Textured Meshes
In most interactive 3D graphics applications, such as video games and virtual/augmented reality applications, the majority of graphics assets are depicted using textured meshes. These meshes are usually designed by graphics artists and creatives, who build complex objects by blending primitive 3D shapes or sculpting complex geometric surfaces.
However, there has also been a significant increase in 3D content captured from real-world scenes in recent years driven by rapid improvements in 3D reconstruction techniques such as structure from motion (SfM) [39, 46, 47] and implicit representations like [34, 9, 42]. While methods exist that can directly generate meshes from input views like [11, 6], most of these methods produce 3D reconstructions in the form of point clouds, voxel grids, or volumetric fields which have to be converted into textured meshes using algorithms like Poisson mapping [19] and marching cubes [30].
When generating a texture map from textured 3D mesh surfaces, a UV parametrization process is performed which involves unwrapping the 3D model by generating a mapping from each point on its surface onto the two-dimensional UV plane. Popular UV mapping techniques include [38, 7, 3]. To reduce distortions that arise during flattening from 3D to 2D, it is common to partition each mesh into multiple patches, whereby the flattening and UV parametrization are done independently on each patch. Multiple ways of decomposing meshes into patches have been proposed [40, 26, 37].
However, noise arising during the 3D reconstruction of geometry often causes unnatural texture seams, inefficient utilization of space on the UV plane, and sub-optimal calculation of the UV mapping. These drawbacks can lead to wasted memory usage while rendering and reduced perceptual quality of the rendered images. These problems are generally exacerbated when the resolution of the texture map is reduced.
2.2 Deep Learning on Meshes
In recent years, many deep learning-based techniques have been proposed for applications related to textured meshes including 3D object recognition [25, 22], detection [5, 21], and segmentation [13, 17]. These efforts have built upon modifications of graph convolutions [20] such as [13, 33, 14]. While graph convolutions are suitable for processing geometry and connectivity of meshes and ordinary Cartesian convolutions are appropriate on 2D texture maps, there are no frameworks for processing them jointly.
A variety of deep models have been proposed for generating 3D meshes from multiple [11, 10] or single images [43, 44]. A key ingredient of these generative models is differentiable rendering, which allows parameters of a 3D scene to be updated using measurements made on rendered 2D views. A plethora of differentiable rendering algorithms have been proposed [18, 29] that implement differentiable rasterization [27, 28] and path-tracing [2, 24].
3 Method
3.1 Overview
GeoScaler takes textured 3D meshes as input, learns an optimal texture downsampling function, and generates optimally downsampled textures. Unlike traditional methods, where texture downsampling is performed without knowledge of the mesh geometry and its UV parametrization (as shown in Eq. 1), GeoScaler exploits these properties in addition to knowledge of the rendering process to generate an optimally downsampled texture map that maximizes the visual fidelity of the rendered views. GeoScaler can be expressed as in Eq. 2, where $\mathbf{T}$ is the original texture map, $\mathbf{T}_{\downarrow s}$ is the texture map downsampled by a factor of $s$, $\mathbf{M}$ is the mesh, and $r(\cdot)$ is the rendering function used to transform the 3D mesh into a 2D view. The mesh M includes the vertex positions $\mathbf{V}$, edge connectivity $\mathbf{E}$, and the UV coordinates of all vertices $\mathbf{U}$. The set $\mathbf{V}=\{v_{j}\}_{j=1}^{N},v_{j}\in\mathbb{R}^{3}$ contains the position of $N$ vertices in the 3D space, $\mathbf{U}=\{(u_{j}^{x},u_{j}^{y})\}_{j=1}^{N},u_{j}^{x},u_{j}^{y}\in\mathbb{R}$ are the UV parameters corresponding to each vertex and $E$ is the adjacency list of connections between vertices:
$\mathbf{T}_{\downarrow s}=\text{DownSampler}_{s}(\mathbf{T})$ (1)
$\mathbf{T}_{\downarrow s}=\text{GeoScaler}_{s}(\mathbf{M}(\mathbf{T},\mathbf{V},\mathbf{E},\mathbf{U}),r(\cdot))$ (2)
GeoScaler computes an optimal downsampling function by training a mesh-specific DNN to downsample its texture map. For every input textured mesh, GeoScaler optimizes the parameters of a deep neural network that downsamples the textures of the mesh in an iterative self-supervised manner, by minimizing the differences between images rendered from the original mesh+ texture and the mesh with downsampled textures. This process can be formulated as shown below in Eq. 3 and Eq. 4 .
$\mathbf{T}_{\downarrow s},\mathbf{U^{\prime}}=\text{DNN}(\mathbf{M}(\mathbf{T},\mathbf{V},\mathbf{E},\mathbf{U}),\theta_{opt})$ (3)
where
$\begin{gathered}\theta_{opt}=\arg\min_{\theta}\text{RenderLoss}(\mathbf{M}(\mathbf{T}_{\downarrow s},\mathbf{V},\mathbf{E},\mathbf{U^{\prime}}),\\ \mathbf{M}(\mathbf{T},\mathbf{V},\mathbf{E},\mathbf{U})).\end{gathered}$ (4)
3.2 Model Architecture
Our proposed deep network architecture enables leveraging cues from the geometric layout of the mesh and its UV parametrization to improve the downsampling performance. Given an input mesh $\mathbf{M}(\mathbf{T},\mathbf{V},\mathbf{E},\mathbf{U})$ with a texture size $H\times W$, the first step involves projecting the three-dimensional geometric properties of the mesh, specifically the vertex positions and normals, onto a two-dimensional map that aligns with the texture map. This can be done by utilizing barycentric coordinates to first smoothly interpolate the geometric features over the surface of the mesh, then using the UV parametrization to flatten them onto a two-dimensional feature map. Since the same UV parametrization is used to project textures onto the surface of the mesh, the texture map will align with the geometric feature maps. The conversion of geometric information to the 2D domain allows it to be filtered with ease via regular convolutional layers. This idea of converting geometric information to texture-aligned 2D features has been previously successfully applied for 3D geometry editing [1]. These feature maps, which contain rich information regarding the three-dimensional properties of the mesh, are concatenated with the texture map before being input to the DNN. This simple step allows the DNN to access the geometric properties of the mesh when determining the optimal parameters for downsampling in a geometry-aware manner.
The input texture map and the geometric features are then processed via an encoder-like module comprising of convolution and pooling layers which extract a set of deep features $(F_{e})$ at the desired reduced resolution of $H/s\times W/s$ and $C_{1}$ channels. This is followed by the GeoCoding module, which explicitly filters the deep features in accordance with the connectivity and layout of the mesh. Prior to reconstructing an RGB texture map, the UVWarper module estimates the warping function to be applied to the texture map and the UV parameters of the mesh vertices. The warping aims to make the texture layout and the UV mapping of the mesh more efficient. Each of these modules is explained in detail in the following sections.
3.3 GeoCoding Module
Due to the limited receptive field of convolution operations, pixels on the boundaries of a texture map are not filtered with their true neighbors on the 3D surface. The GeoCoding (short for ”geometric encoding”) module aims to address artifacts that arise due to these discontinuities of the texture map. The structure of the GeoCoding module is depicted in Fig. 3.
Figure 3: Overview of the GeoCoding module. Features obtained from the texture encoder are mapped to vertices on the surface of the mesh using the UV parametrization of the mesh. Performing graph convolution operations on these mapped features allows combining geometric information from the mesh with texture information, which reduces artifacts at texture map discontinuities. The output features are then interpolated using Barycentric coordinates and reprojected back to the UV plane.
The first step involves using the UV parametrization $(\mathbf{U})$ to sample a $C_{1}$-dimensional feature embedding ($f_{i}$) at the location of each of the $N$ mesh vertices from $F_{e}$, the $C_{1}\times H/s\times W/s$ dimensional feature map produced by the CNN-based encoder. Eq. 5 denotes this process:
$\begin{gathered}u_{i}^{x},u_{i}^{y}=\textbf{U}(v_{i}),i=1...N\\ f_{i}=F_{e}[u_{i}^{x},u_{i}^{y}]\end{gathered}$ (5)
If a vertex appears more than once in the UV parametrization, we average the samples from each position. After sampling, we obtain a graph with $N$ vertices and a feature embedding $(f_{i}\in\mathbb{R}^{C_{1}})$ associated with each vertex. Unlike the 2D feature map which suffers from discontinuities due to the UV mapping, these vertex-wise feature embeddings $(f_{i})$, albeit sparse, lie adjacent to their true neighbors on the 3D surface and are connected to them via the edges of the mesh.
The vertex-wise feature embeddings are filtered using residual blocks built using simple graph convolution operations as shown in Eq. 6 [35].
$\begin{gathered}f^{\prime}_{i}=\theta_{1}f_{i}+\theta_{2}\sum_{j\in N(i)}e_{ij}f_{j},\end{gathered}$ (6)
where $N(i)$ is the set of neighbors of $i$, $e_{ij}$ is the weight between nodes $i$ and $j$, and $\theta_{1}$ and $\theta_{2}$ are learned parameters. We designate the edge weights $e_{ij}$ to equal the Euclidean distance between the mesh vertices. Performing the convolutions in the geometric domain ensures that features that are proximal in the 3D domain on the surface of the mesh, but lying on distant seams of texture maps, are appropriately fused together. The sequence of graph convolutions outputs a set of vertex-wise feature vectors of length $C_{2}$.
To reconstruct a two-dimensional map from geometrically filtered features, we first smoothly interpolate the features sampled at each vertex across the faces using barycentric coordinates. Next, the UV parametrization is again used to project features on the mesh surface onto the 2D UV plane. The output of the GeoCoding module, a two-dimensional feature map $(F_{g})$ of size $C_{2}\times H/s\times W/s$ is generated. The obtained feature set $F_{g}$ is compensated for geometric discontinuities and distortions present in the texture map. If vertices are extremely sparse, and the sampled vectors at the vertices fail to retain sufficient information, $F_{g}$ is added with a skip residual from $F_{e}$.
3.4 UVWarper Module
[…]
Finally, the downsampled RGB texture map $\mathbf{T}_{\downarrow s}$ is reconstructed from $F_{w}$. This downsampled texture map $\mathbf{T}_{\downarrow s}$ is overlaid on the surface of the mesh using the offset-adjusted UV parameters U’, yielding the mesh representation $\mathbf{M}(\mathbf{T}_{\downarrow s},\mathbf{V},\mathbf{E},\mathbf{U^{\prime}})$ as the final output.
3.5 Rendering Loss and Optimization
To optimize the parameters of the DNN, the differences between the rendered views generated from the original mesh and the output mesh are minimized. At each iteration of training, a batch $\mathbf{P_{K}}=\{P_{j}\}_{j=1}^{K}$ of $K$ camera poses ($P_{j}$) focusing on the mesh are sampled randomly from a predefined range $\mathbb{S}$ of plausible and reasonable viewing conditions. The range $\mathbb{S}$ is selected such that all reasonable viewing angles and distances are included within it. At each iteration of training, $K$ images are rendered from both the original mesh and the processed mesh with the cameras placed and oriented according to the set of sampled poses $\mathbf{P_{K}}$ as shown in Eq. 7
$\begin{gathered}\mathbf{I_{K}^{orig}}=\{r(\mathbf{M}(\mathbf{T},\mathbf{V},\mathbf{E},\mathbf{U}),P_{j})\}_{P_{j}\in\mathbf{P_{K}}}\\ \mathbf{I_{K}^{ds}}=\{r(\mathbf{M}(\mathbf{T_{\downarrow s}},\mathbf{V},\mathbf{E},\mathbf{U^{\prime}}),P_{j})\}_{P_{j}\in\mathbf{P_{K}}}\end{gathered}$ (7)
where $r(\mathbf{M}(\cdot),P)$ is the rendering function that projects a textured mesh $\mathbf{M}(\cdot)$ to an RGB image from a viewing pose $P$, and $\mathbf{I_{K}^{orig}}$ and $\mathbf{I_{K}^{ds}}$ are the sets of images rendered from the original mesh and the mesh with the downsampled texture, respectively. From these rendered images, a combination of 2D loss functions including L2 loss, SSIM [45] loss, and VGG-19 [41] based perceptual loss [15] are computed. The losses of each pair of rendered images are summed to obtain the rendering loss for that iteration as shown in Eq. 8.
$\begin{gathered}\text{RenderLoss}=\sum_{j=1}^{K}[\alpha_{1}\mathcal{L}_{L2}(I_{j}^{orig},I_{j}^{ds})\\ +\alpha_{2}\mathcal{L}_{SSIM}(I_{j}^{orig},I_{j}^{ds})+\alpha_{3}\mathcal{L}_{perceptual}(I_{j}^{orig},I_{j}^{ds})]\end{gathered}$ (8)
Given the availability of a plethora of differentiable rendering algorithms, one can be chosen that closely matches the target application. The performance of the model is validated after every few iterations by averaging the losses computed from images rendered from a set of $Q$ of poses sampled uniformly from $\mathbf{P_{K}}$.
4 Experiments
Figure 4: The newly proposed 3DSet5 dataset. Meshes in the 3DSet5 dataset are unprocessed and unrefined after reconstructing from multiple images of the objects and contain numerous geometric and textural irregularities.
Figure 5: Qualitative results comparing GeoScaler with bicubic downsampling for 4x and 8x downsampling. The top row of meshes is from Google’s Scanned Object Dataset, the middle row is from our 3DSet5 mesh dataset, and the bottom row is from the Textured Mesh Quality Assessment dataset. For each mesh result, the top row is the Ground Truth texture, the middle row shows results from GeoScaler, and the bottom row shows results using Bicubic.
Scale
(Texture Size)
4x
8x
Dataset Mesh / Method
GeoScaler
PSNR(dB) / SSIM
Bicubic
PSNR(dB) / SSIM
GeoScaler
PSNR(dB) / SSIM
Bicubic
PSNR(dB) / SSIM
3DSet5 avocado 35.04 / 0.9559 32.48 / 0.9336 33.13 / 0.9309 29.12 / 0.8868
banjoman 41.92 / 0.9855 40.81 / 0.9823 35.18 / 0.9481 33.77 / 0.9389
bevo 30.06 / 0.9300 27.94 / 0.8982 28.68 / 0.8990 25.87 / 0.8465
clothdolls 35.96 / 0.9650 33.53 / 0.9440 32.46 / 0.9269 28.22 / 0.8580
stew 32.28 / 0.9457 29.58 / 0.9119 29.78 / 0.9041 26.17 / 0.8431
GSO shoe 28.12 / 0.9294 26.33 / 0.9155 26.07 / 0.9131 23.94 / 0.8760
shoe2 31.63 / 0.9601 30.11 / 0.9508 26.74 / 0.9118 25.28 / 0.9041
spiderman 36.53 / 0.9820 35.35 / 0.9785 33.44 / 0.9683 31.79 / 0.9606
TMQA mask 32.70 / 0.9516 31.50 / 0.9394 26.85 / 0.8631 25.46 / 0.8393
pizza 33.57 / 0.9122 32.58 / 0.8874 30.13 / 0.8306 28.54 / 0.8023
Table 1: Quantitative results comparing GeoScaler with bicubic downsampling for 4x and 8x downsampling. Each cell contains the PSNR(dB) and SSIM value. GeoScaler outperformed bicubic downsampling by a significant margin across all datasets.
Scale
GeoScaler
w/o UVWarping
w/o UVWarping &
w/o GeoCoding
4x
35.05 /
0.9564
34.45 /
0.9515
34.18 /
0.9486
8x
31.85 /
0.9218
30.54
0.8983
30.15 /
0.8895
Table 2: Ablation results of the main modules in the GeoScaler architecture. Average PSNR(dB)/SSIM obtained on the 3DSet5 dataset before and after ablating the GeoCoding and UVWarp modules.
Figure 6: Trade-off between rendering memory consumption and visual quality. Textures produced by GeoScalar produce superior visual quality while consuming the same rendering memory compared to textures produced by bicubic downsampling.
4.1 Implementation Details
Our DNN model contains 4 residual blocks in the downsampling encoder and 2 graph residual blocks in the GeoCoding module. We chose $C_{1}=C_{2}=64$, bringing the total number of parameters in the network to 0.57M. To compute the rendering loss, we used the differentiable renderer proposed in SoftRasterizer [27], along with a Soft Phong Shader [4] whose implementations are available in PyTorch3D [16]. At each iteration, we rendered $K=4$ views when computing the loss. For the rendering step, we uniformly illuminated the scene with a white ambient light. The AdamW optimizer [32] was employed for the training process. We used a peak learning rate of $35e-$ 3 with a cosine annealing schedule [31] and a warm-up over 1000 iterations. We found that it is sufficient to train the DNN over 15000 iterations for convergence. When validating the model, we used $Q=60$ poses sampled uniformly over the range of camera poses.
4.2 Datasets
We tested GeoScaler on a wide variety of textured meshes selected from multiple datasets including Google’s Scanned Object (GSO) dataset [8] and the Textured Mesh Quality Assessment (TMQA) dataset [36].
We also studied the performance of Geoscaler on unprocessed and unrefined textured meshes which are directly generated from multiple images of real-world 3D objects, which contain geometric artifacts and irregularities. To that end, we share a small dataset of five textured meshes that did not undergo any post-processing or refinement and present the performance of GeoScaler on it. We dub this dataset 3DSet5 and its meshes are shown in Fig. 4. The meshes in the 3DSet5 were generated from multiple input views (approx. 50) using Agisoft’s Metashape application. The meshes in 3DSet5 contain approximately 10000 vertices and 20000 faces and have texture maps of size 2048x2048.
4.3 Results and Ablation Study
We measured the quality gains provided by GeoScaler in terms of peak signal-to-noise ratio (PSNR) and the Structural Similarity Index Metric (SSIM) as shown in Tab. 1. GeoScaler was able to significantly outperform existing methods by improving downsampling quality at both 4x and 8x scale factors. In some cases, textures obtained when performing 8x downsampling using GeoScaler produce better results than even textures obtained by 4x bicubic downsampling. This is evident on meshes avocado, bevo, and stew from the 3DSet5 dataset, as shown in Fig. 5. the objects shoe and spiderman in Fig. 5 are from Google’s Scanned Objects dataset [8], mask and pizza are from the Textured Mesh Quality Assessment dataset [36], and bevo and stew are from 3DSet5. It can be noticed that GeoScaler displays significantly superior retention of details and fine structures contained in the original texture. GeoScaler also prevents artifacts at seams on textures, as shown in 8x downsampling results of bevo (horizontal artifact in the left column of 8x bicubic result).
Next, we ablated the UVWarping module and the GeoCoding module to validate their importance. Tab. 2 shows the average PSNR and SSIM values over the 3DSet5 dataset when the modules are ablated. Detailed visual results are available in the supplementary materials. In supplementary material, we also show that textures generated by GeoScaler provide perceptually better results than textures generated natively at lower resolution using MetaShape over the 3DSet5 dataset.
Using GeoScaler, the trade-off between memory usage while rendering and visual quality is greatly improved, with noticeably improved visual quality of the rendered contents even though much lower rendering run-time memory was required. The average reduction in SSIM is plotted against memory consumed while rendering in Fig. 6 to highlight this aspect. The memory consumed is equal to the total number of pixels in the texture, multiplied by its bit-depth and by the number of color channels.
5 Conclusions
We presented a method of downsampling texture maps that is superior to existing methods using a unified representation of texture maps and the geometries of 3D meshes that they are painted on. This framework is also applicable for downsampling other feature maps associated with meshes, such as roughness and specularity.
GeoScaler-based downsampling of textures is intended to enable the deployment of compact graphical assets in applications for devices having low computation and memory budgets such as smartphones and VR headsets while maintaining high visual quality. It also facilitates remastering high-quality graphical assets for video games on devices with lower rendering capabilities.
Supplementary Material
6 More Results
Visual results of meshes (avocado, banjoman, and shoe2) included in the full paper are shown in Fig. 7 for the sake of completeness. All mesh .obj files with the original, bicubic downsampled, and GeoScaler downsampled texture maps are also available in the submitted supplementary material.
Figure 7: Qualitative results comparing GeoScaler with bicubic downsampling for 4x and 8x downsampling. For each mesh result, the top row is the Ground Truth texture, the middle row shows results from GeoScaler, and the bottom row shows results using Bicubic.
A more detailed mesh-wise ablation is shown in Fig. 8 and Fig. 9 for 4x and 8x downsampling respectively. Fig. 8 and Fig. 9 also show the quality of images rendered using textures generated at lower resolutions by Metashape directly while performing the 3D reconstruction process. These results suggest that downsampling textures after reconstructing 3D scenes at higher resolutions (using tools such as MetaShape) can provide higher-quality renders than generating low-resolution textures directly. In these tables, it is also evident that Lanczos resampling and bicubic resampling are extremely similar quantitatively. Given this observation and the widespread popularity of bicubic resampling, the results of Lanczos resampling were omitted from the full paper for the sake of brevity. Qualitative results of the ablations for stew and bevo are also shown in Fig. 11.
Figure 8: Mesh-wise Ablations for 4x Downsampling. GeoScaler outperforms traditional methods and also performs better than generating textures natively at lower resolutions
Figure 9: Mesh-wise Ablations for 8x Downsampling. GeoScaler outperforms traditional methods and also performs better than generating textures natively at lower resolutions
Figure 10: Comparing Texture maps generated by bicubic downsampling and GeoScaler. Textures generated by GeoScaler have larger areas assigned to regions with high detail density (as shown by red boxes) and are more sharpened to compensate for the downsampling
Figure 11: Visual Results of Ablation Study. The independent contributions of the GeoCoding Module and the UVWarping module are illustrated. GeoCoding stitches artifacts arising due to seams in texture maps while UVWarping improves the details in the rendered images
7 A Closer Look at the Generated Textures
Instead of analyzing rendered images, we study the texture maps themselves. The difference between texture maps downsampled by bicubic and GeoScaler is depicted in Fig. 10, which contains crops of 8x downsampled texture maps of avocado, clothdolls and stew meshes. It can be observed that the texture maps generated by GeoScaler look over-sharpened. However, it is important to note that texture maps are first wrapped onto surfaces of meshes (whose surface areas remain unchanged throughout the downsampling process) before views are rendered using them. As surface areas of meshes remain unchanged, the downsampled textures have to be stretched/upsampled to cover them. The over-sharpening and the upsampling compensate for each other’s detrimental distortion and produce a rendered result that resembles that of the original texture. GeoScaler is capable of estimating the degrees of (over-)sharpening needed to apply on different regions of texture maps based on ratios of areas of faces on mesh surfaces and their corresponding UV planes.
In Fig. 10 the effect of the UVWarper module is also visible. The regions marked by the red boxes occupy a larger area in texture maps produced by GeoScaler compared to the area they cover in the texture maps produced using bicubic downsampling. Without any additional semantic guidance/supervision, the UVWarper is capable of recognizing areas of important textural information and consequently assigning larger areas to them in the generated texture map. The warping of texture maps reduces the loss of information occurring during the downsampling of the textures and enhances the perceptual quality of rendered images.